# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

### v2

@@ 3 Results
We evaluated our workflow using experimental datasets associated with meaningful scientific context. SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data. From here, we selected three experimental datasets from astronomical observations (Bode’s Law, Hubble’s Law, and Kepler’s Law). To these, we added two experimental chemistry datasets: Langmuir and Dual-site Langmuir adsorption isotherm models [9].
Figure 18 in SI illustrates all the datasets with the target expression for each. We also tested our workflow using a dataset on friction losses in pipe flow [89]. This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6]. We included the whole dataset for each problem in the prompt to LLMs at the start and in each iteration. Because the Hubble’s dataset has an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to LLMs but used the original dataset when optimizing constants in SciPy. This reduced the tokens (and cost) of running the models. We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
We conducted eight sets of experiments for each model (GPT-4 and GPT-4o) on six datasets, running each set five times to evaluate its robustness. In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data. Empirical relations often involve field-specific operators, so the equation search should account for this. We refer to this set of tests as an “easy search" because the space of possible equations is constrained to that generated by the operators in our dataset. In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search. All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on. In the following text, we discuss each dataset individually and analyze their performance. The score at each iteration is a total of x/5 independent runs where target model expressions were found.
@@
Temperature is a hyperparameter used in stochastic models like LLMs to regulate the randomness of the model output [90, 91]. It adjusts the probabilities of the predicted words in the softmax output layer of the model. Lowering temperature favors words with higher probability, so when the model randomly samples the next word from the probability distribution, it will be more likely to choose a more predictable response. We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the datasets.
Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which was often a first guess in the GPT-4 model when few operators are made available. However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or context inhibited performance (Figure 7). This dataset is particularly noisy, and including the noisy data as context actually inhibits the search as observed by consistent scores of 5/5 from both models when they were run without the data. Furthermore, the scratchpad in GPT-4o reveals that when the model first guesses Hubble’s Law, it then tends to reason and deduce a more complex, non-linear model in context of the noisy data.
Figure 7: GPT-4 and GPT-4o results on Hubble’s data. In the GPT-4 model, all the trends except ‘Without Context’ overlap at score 5 for the “Easy Search” plot, and the trends ‘With All Tools’ and ‘Without Data’ overlap at score 4 for the “Hard Search”. In the GPT-4o model, the trends for ‘Without Data’ and ‘Without Context’ overlap at score 0 for the “Easy Search” plot, and all the trends except ‘Without Data’ overlap at score 0 for the “Hard Search”. Transparency and different marker styles are used to differentiate these trends.
Bode: The results for Bode’s Law indicate that similar to other SR programs [88], LLMs encounter difficulty in accurately rediscovering the target model expression. Curiously, it performs best when the context is excluded and worst when the data is excluded. This suggests that our context was counterproductive for the search, and purely reasoning about the data led to greater success, with GPT-4o leading the performance (Figure 8). Scratchpad shows that GPT-4o refers to Bode in a few experiments, specifically in relation to the ’semi-major axis’ variable identity provided in the context, which may have been overly specific. However, it never suggested the target model expression in those instances. We also notice GPT-4o rediscovering the target model expression in the first few iterations from just the data patterns (Figure 9). However, both GPT-4 and GPT-4o find the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
Figure 8: GPT-4 and GPT-4o results on Bode’s data. For the “Hard Search” in the GPT-4 model, all the trends except ‘Without Context’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
Figure 9: GPT-4o reasoning on Bode’s data. The scratchpad snippet is from the first iteration in Run 4 with all tools on for hard search.
Kepler: We find strong evidence that the prompt has triggered LLM’s memorization of Kepler’s Law: the scratchpad reveals that both GPT-4 and GPT-4o models associate the variable names and data in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification. Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in OpenAI’s training data than the other relationships. However, scratchpad also shows that although GPT-4o infers Kepler’s relationship from data and/or variable identity, it does not always suggest the target model expression but reasons with alternative expressions based on dataset patterns and feedback from prior suggestions. Both the models also often find the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$. However, SciPy optimizes and fits $c_{2}$ with a floating point power ($\approx 1.5$). We did not consider this a rediscovery of the target model as such expressions are not dimensionally consistent.
Figure 10: GPT-4 and GPT-4o results on Kepler’s Third Law data. For the “Hard Search” in the GPT-4o model, the trends for ‘Without Data’ and ‘With All Tools’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round. In easier searches, both GPT-4 and GPT-4o consistently found it within 15 iterations; inclusion of context and data seemed to improve performance. However, the hard search was much more challenging, with Langmuir discovered a few times, mainly with all tools on. An interesting observation about the GPT-4o model is that it references Langmuir and other adsorption models (Freundlich or BET), particularly for experiments with easy search and without data, where the model focuses more on the adsorption context. However, similar to other benchmark datasets, it does not always propose the target model expression; instead, it methodically analyzes the information it receives, including the context and feedback.
Figure 11: GPT-4 and GPT-4o results on Langmuir’s adsorption data. For the “Hard Search” in the GPT-4 model, all the trends except ‘With All Tools’ overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions. In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 12). However, extrapolation shows this model is not theoretically correct (Figure 19 in SI). This may also explain why we see a decreasing score trend, even after identifying the target expression, compared to other datasets (Figure 13). We analyzed the LLM-generated expressions from both dual-site and Langmuir adsorption datasets by checking whether they satisfy thermodynamic constraints by running through a computer algebra system built in our previous work [28]. We consistently found one more accurate expression at the expense of an added complexity for Henry’s Law term, which models adsorption at low pressures well.
Figure 12: Pareto fronts for dual-site Langmuir dataset. The black line represents the best total front from the five independent runs. The target model is labeled as a blue star.
Dual-site Langmuir dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92]. Because the target model expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search in GPT-4 model. We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop). We found that when using basic operators (easy search) and all tools on, GPT-4 was getting close to finding the target model (Figure 12) upon running 15 iterations with this feedback loop. To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed. Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR.
Figure 13: GPT-4 and GPT-4o on dual-site Langmuir’s adsorption data. All the trends in GPT-4 model overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
@@
However, the same prompts were more successful with GPT-4o model, which reasoned better with data and context. For experiments without data, GPT-4o relied on the context of general adsorption behavior. It initiated with expressions based on Langmuir, Freundlich, and occasionally Temkin and BET theories. Interestingly, we observed the highest score on this dataset when no context was provided, and the target model was rediscovered through dataset patterns and iteration feedback. Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in the datasets related to saturation behavior and leading to modifications of the Langmuir expression to derive the dual-site behavior.
@@
Figure 14: Instances of GPT-4o reasoning on dual-site Langmuir dataset. The top snippet of the scratchpad is from Run 1, and the bottom is from Run 2. Both are for runs with all tools on and iteration 4.
@@
Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to LLMs exceeded the token limit. Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations. So, we developed a cost-saving scheme: only send a portion of the data in the prompt to LLM, while fitting and evaluating the generated expressions using the whole dataset (Figure 20 in SI).
@@
Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging. We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions. So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error. Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to LLMs for feedback. To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
@@
Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6]. We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better. We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) data points and the second (S2) with another 36 (20%) data points. Overall, feeding in more data generated better-fitted and more complex expressions. Figure 15 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^). We also list the complexities and MAEs for these expressions in Table 1 in SI.
@@
Figure 15: Models for Nikuradse dataset. Here, P refers to prompt versions and S refers to dataset. The expressions generated by the GPT-4o model are marked with an ‘o’ at the end. We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig. 15 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways. However, GPT-4o model only shows unphysical behavior in prompt version 1 with fewer data. Both the models generalize better with fewer data in prompt version 3 where we challenge by providing information about the MAE and complexity of the BMS model.
@@
We found optimal expressions from GPT-4 and GPT-4o with a complexity of 41 and 27 and an MAE of 0.01086 and 0.00924, respectively (Figure 15 C). GPT-4o performs qualitatively better, generating a more accurate expression at a significantly lower complexity than GPT-4. The MAE from both models is at least two to three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS). BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392. However unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression. Our model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions. The optimal expression chosen is the best from 5 runs (1/153*5 or 1/765). The incremental suggestions from LLMs for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS. However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost. For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
@@
BMS samples expressions from a probability distribution. Therefore, running for a long time is expected to be characterized by equilibration, in which the expressions converge to a region in the stationary distribution of expressions, after which no significant improvement would be observed apart from continuing to explore the region of most likely expressions. Genetic algorithms like Eureqa [3] do not converge to a stationary distribution and nor does our approach. On the contrary, running a chatbot for a very long time leads to degradation of its answers as the context gets longer [76, 77], so, in principle, our method may exhibit similar degradation with very long runs. However, we expect this effect to be mitigated because we discard the majority of the context after each iteration, only passing the top and recent expressions.
We also show a model from Eureqa in [6] which was run with default operator penalties and selected for the best expression from at least $10^{13}$ ones. The result from the EFS [85] model (best fit selected from 100 runs) in [6] is comparable to our model with an MAE of 0.00941. EFS is based on sparse regression that uses a genetic algorithm to generate basis functions automatically. EFS is known to be highly efficient, delivering expressions in seconds.

### v3

@@ 3 Results
We focused on applying LLM-based SR to experimental datasets, exploring its potential to uncover interpretable scientific relationships rather than benchmarking on synthetic datasets. Specifically, we investigated two well-known chemistry datasets: Langmuir and dual-site Langmuir adsorption isotherm models [60]. These two datasets were also examined in our previous work [23, 33], which enables us to directly compare this LLM-based SR approach with traditional SR methods. We further tested our workflow on a challenging dataset of friction losses in pipe flow [78], which does not have an established target model and has been the subject of prior studies by other SR programs [40, 13]. We used Nikuradse data from [40], which was sourced directly from the original reported literature [78], without any modifications.
We also tried datasets corresponding to Hubble’s Law, Kepler’s Law, and Bode’s Law (from PySR’s empirical benchmark [25]). However, these were highly susceptible to different forms of leakage, and we therefore include those results only in the SI section 5.2 for interested readers. The scientific context was often sufficient for the LLMs to infer the correct law name and the target model expression in the first iterations. Additionally, the data, such as the orbital periods of the planets in the solar system, were recognizable to the LLMs even without the scientific context, as it appears in the Wikipedia article on Kepler’s Law [21]. Figure 22 in SI illustrates all the explored datasets with their respective target model expressions.
3.1 Experimental settings
We conducted experiments using OpenAI’s GPT-4 and GPT-4o models in ‘Easy Search’ and ‘Hard Search’ settings. ‘Easy Search’ includes only the basic binary operators (++, -, *, $\div$). In addition to this, the ‘Hard Search’ includes standard unary operators (sqrt, log, exp, square, cube) to evaluate a more challenging search space. Additionally, we compared the performance of LLMs on following mathematical constraints (derived from background theory of data) with our previously explored work using computer algebra systems [33].
Our workflow has three main components:
  • Data: sent as input in natural language. For instance, Langmuir’s data in the prompt looks like: “Your job is to find 3 expressions that approximately describe the dataset with dependent variable, y: [33, …] and independent variable, $x_{1}$: [34, …]”
  • Context: is sent as a single-line text input about data and its variable identities; for instance, the context used for Langmuir’s data is: The data is about nitrogen adsorbing onto mica, where the independent variable (x1) is pressure, and the dependent variable (y) is loading.
  • Scratchpad: frames the task as a reasoning problem in the context of data and free-form background information. For example, a scratchpad output from the LLMs for Langmuir’s data looks like: Looking at the dataset, we can see that ‘y decreases as $x_{1}$’ increases. Therefore, it seems to have an inverse relationship. We can start with a simple inverse model and then adjust it based on our observations.
@@ 3.1 Experimental settings
Temperature is a dimensionless hyperparameter used in stochastic models such as large language models (LLMs) to regulate the randomness of output generation [2, 1]. It adjusts the probabilities of the predicted words in the softmax output layer of the model. Lowering the temperature favors words with higher probability, so when the model randomly samples the next word from the probability distribution, it will be more likely to choose a more predictable response. We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the experiments.
Henceforth, the results are divided into four parts –
  • Section 3.2 presents the results from Langmuir’s adsorption data, using Pareto fronts and ablation studies on each experimental factor.
  • Section 3.3 provides analogous results for the dual-site Langmuir dataset.
  • Section 3.4 compares the quality of LLM-generated expressions with constraints added from background knowledge on both Langmuir and dual-site Langmuir datasets.
  • Section 3.5 demonstrates the applicability of this novel method on a larger dataset through the Nikuradse dataset [78], which is a multivariate experimental dataset that correlates friction with pipe roughness and Reynolds’ number, comprising over 350 measurements.
3.2 Langmuir’s adsorption model
The Pareto fronts below illustrate the overall performance of the GPT-4 and GPT-4o models in both ‘Easy Search’ and ‘Hard Search’ settings. Each experiment was conducted five times, and the expressions from these five runs are represented as scatter points in five different colors. Each Pareto front is also shown in its respective color, with an overall Pareto front indicated by black dashed lines. The overall front is created by pooling the results from all five Pareto fronts to present a single representation of the best performance across all runs. No model is less complex or has less MSE than the one model in the Pareto frontier. The target model expression is marked with a blue star, and its corresponding complexity and mean squared error (MSE) are indicated by grey dashed lines across the axes. As shown in Figure 8, the target model expression for Langmuir was identified in all cases, except for the ‘Hard Search’ with GPT-4.
Figure 8: Pareto fronts for Langmuir’s adsorption model. The target model is expressed as $\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}$ with an MSE of 0.12263 and complexity of 7 (blue star).
We also performed ablation studies under four configurations: without context, without data, without scratchpad, and with all tools enabled (the latter corresponding to the Pareto Fronts shown in 8). Figure 9 analyzes the performance of both models at each iteration level for all experimental settings.
Figure 9: Ablation studies on Langmuir’s adsorption data. The score at each iteration represents the number of runs (out of five) in which the models successfully recovered the target model expression.
Langmuir’s model is obscure and was hard to discover in the first rounds when no context was provided. Without data, the models rely more on the adsorption context, which increases the likelihood of accurately inferring the target model expression early on. However, while the models reference Langmuir, they also refer to other popular adsorption models like Freundlich or BET, and do not always propose the target model expression without considering the dataset pattern (Figure 10).
Figure 10: Instances of GPT-4o reasoning on Langmuir dataset. The snippet of the scratchpad is from an experiment in Easy Search and without data (run #1\#1, iteration #1\#1). The purple-highlighted text shows references to popular adsorption models, the yellow ones show general observations about context and data, and the blue-highlighted one shows the generated target model equation.
3.3 Dual-site Langmuir’s adsorption model
Figure 11: Pareto fronts for dual-site Langmuir’s adsorption model. The target model is expressed as $\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}+\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}$ with an MSE of 0.00079844 and complexity of 15 (blue star).
Figure 12: Ablation studies on dual-site Langmuir’s adsorption data. The score at each iteration represents the number of runs (out of five) in which the models successfully recovered the target model expression.
Figure 13: Extrapolation reveals theoretical
inconsistencies of models.
Dual-site Langmuir’s model dataset is particularly challenging for SR because the target model does not significantly fit the data much better than many other shorter expressions. This was also observed in previous literature studies [23, 33] using three SR algorithms: Bayesian-based SR (BMS) [40], genetic programming-based SR (PySR) [25], and mixed-integer nonlinear programming-based SR [4]. As seen in Figure 11, the models generate other expressions with similar accuracy and half the complexity that reach almost similar accuracy as the target model expressions, creating a plateau from complexity 7 onward. However, most of these shorter expressions are inconsistent with the theory. For instance, extrapolation of such an expression (with complexity 11) generates a negative value for adsorption loading at lower pressures (Figure 13).
Our prompts instruct LLMs to generate simple expressions, so they output shorter equations with reasonable MSEs; parsimony is a common goal in SR. Since the target model expression is longer for this dataset, we extended the run to 50 iterations and modified the feedback loop to pass only the top five expressions ranked by MSE in order to send more accurate and longer expressions in the feedback. This led to a decline in scores over iterations when the model initially discovered the ground-truth expression but later discarded it, as newer expressions with lower MSE replaced it in the top five, as shown in the ablation studies in Figure 12. This was also observed in case of Langmuir’s data (Figure 9) with GPT-4o. However, this does not affect the overall outcome since the best expressions that balance both accuracy and complexity were also evaluated in the Pareto fronts, as shown in Figures 8 and 11.
@@ 3.3 Dual-site Langmuir’s adsorption model
We did not perform experiments with ‘Hard Search’ for this dataset, as the ‘Easy Search’ case already presented significant challenges, especially for GPT-4 model. GPT-4o model was more successful and reasoned better with data, context, and scratchpad. For experiments without data, like in Langmuir’s case, the models relied on the context of general adsorption behavior. It initiated with expressions based on Langmuir, Freundlich, and occasionally Temkin and BET theories. Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in datasets related to saturation behavior, leading to modifications of the Langmuir expression to derive the dual-site behavior.
@@
Figure 14: Instances of GPT-4o reasoning on dual-site Langmuir dataset. The top snippet of the scratchpad is from run 1 (iteration # 5), and the bottom is from run 2 (iteration # 4). Both are for experiments with all tools. The purple highlighted text shows the reference to Langmuir adsorption model, yellow ones show the general observations about context and data pattern, and blue highlighted one shows the generated target model equation.
3.4 Constraining equation search with natural language prompts
SR algorithms traditionally explore the equation search space randomly, which is inefficient for scientific datasets following background knowledge, often leading to inconsistent equations, as seen in the dual-site Langmuir case. In our previous work [33], we used a computer algebra system to limit the search space of two SR algorithms, PySR (genetic algorithm-based) and BMS (Markov chain Monte Carlo-based), by applying symbolic mathematical constraints. This approach allowed us to generate expressions that were meaningful and theoretically consistent. However, incorporating symbolic constraints is challenging in an existing SR package as it requires software development across multiple layers. In this work, we leverage LLMs to simplify this process, and we express these constraints in natural language in the prompts, which is far simpler and reduces human effort. The earlier study, with Fox et. al [33], also examined single-component adsorption datasets, two of which overlap with the datasets explored in this work – Langmuir’s and dual-site Langmuir’s adsorption models. The constraints are formulated as follows:
$\lim_{p\to 0}f(p)=0$ (1)
$\lim_{p\to 0}f^{\prime}(p)<\infty$ (2)
$\forall p>0\qquad f^{\prime}(p)\geq 0$ (3)
Here, Constraint 1 ensures that all molecules must desorb, and loading cannot be negative in the limit of zero pressure. Constraint 2 requires that the slope of the isotherm must be a positive finite constant in the limit of zero pressure. And, Constraint 3 requires that loading does not decrease with increasing pressure (the isotherm is monotonically non-decreasing) for all ($\forall$) positive values of pressure. We constructed the same constraints in our prompts:
  1. 1.
    Passes through origin: as $x1\rightarrow 0^{+}$, $y(x1)\rightarrow 0$.
  2. 2.
    Finite, positive initial slope: as $x1\rightarrow 0^{+}$, $0<dy/dx_{1}<\infty$
  3. 3.
    Monotonic non-decrease: for all $x1>0,dy/dx1\geq 0$
We demonstrate the effect of adding constraints on the generated expressions in the GPT-4o model in comparison with PySR and BMS in Figure 15.
(a) Effect of constraints on Langmuir dataset.
(b) Effect of constraints on the dual-site Langmuir dataset.
Figure 15: Comparison of generated expressions with constraints on and off conditions across GPT-4o, PySR, and BMS methods. The data for PySR and BMS is as reported in [33]. The data for GPT-4o is from the ‘Easy Search’ experiments to align with the operators used in both PySR and BMS, as well as with all tools. For each run, the passing rate was calculated as the proportion of expressions (across all iterations within that run) satisfying each constraint. Error bars represent the standard error of the mean across the five runs.
Notably, the largest improvements are observed with Constraint 1 across both datasets (see Figures 15 and 16). Despite its simplicity, the training data may not accurately reflect the underlying thermodynamic principles. An expression that fits the data well but does not approach zero loading during interpolation is unphysical and should not appear on the Pareto front as a top-performing one. The effect becomes more noticeable for the dual-site Langmuir case, and the benefit of constraining the search space becomes clearer, as the target model achieves accuracy comparable to shorter expressions that violate domain-specific constraints (see Figure 13).
(a) Distribution of root mean square error (RMSE) for Langmuir expressions.
(b) Distribution of root mean square error (RMSE) for dual-site Langmuir expressions.
Figure 16: Distribution of generated expressions with constraints on and off conditions across the GPT-4o model. Green areas represent models that passed constraint C1, while orange areas represent models that failed across the five iterations. The dashed line indicates the RMSE of the target model, and the y-axis shows the percentage of total expressions in each category.
3.5 Performance of LLM-SR on open scientific problems
@@
Finally, to test the scalability of the approach to larger and multivariate datasets, and to evaluate a problem without a known target model in literature, we examined the Nikuradse dataset, which is experimental data on turbulent friction in rough pipes by Johann Nikuradse in the early 1930s [78]. The data correlates the friction factor ($\lambda$) with relative roughness (k/D) and Reynolds’ number (Re), and contains over 350 measurements. Including the entire dataset in our prompts to LLMs exceeded the token limit, and even with long context windows, analyzing large datasets would be expensive, as each iteration is more costly, and generating longer expressions requires more iterations. Therefore, we developed a cost-saving scheme: only send a portion of the data in the prompt to the LLM, while fitting and evaluating the generated expressions using the entire dataset (Figure 26).
@@
Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimization was also more challenging. We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions. So, we optimized the constants ten times for each expression, then selected the one with the lowest mean absolute error (MAE). We evaluate performance on this dataset computing for MAE to align with the performance metrics used in the literature by other SR programs. Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to LLMs for feedback. To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
@@
Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs discussed by Guimerà, et al [40]. We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing the MAE achieved by a literature model (without leaking the model), and challenging it to do better. We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) data points and the second (S2) with another 36 (20%) data points. Overall, feeding in more data generated better-fitted and more complex expressions. Figure 17 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators (++, -, $\div$, $\times$ and ˆ). We also list the complexities and MAEs for these expressions in Table 1.
@@
Figure 17: Models for Nikuradse dataset. Here, P refers to prompt versions and S refers to the dataset. The expressions generated by the GPT-4o model are marked with an ‘o’ at the end. We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig. 17 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways. However, GPT-4o model only shows unphysical behavior in prompt version 1 with fewer data. Both the models generalize better with fewer data in prompt version 3, where we challenge by providing information about the MAE and complexity of the BMS model.
@@
We found optimal expressions from GPT-4 and GPT-4o with a complexity of 41 and 27 and an MAE of 0.01086 and 0.00924, respectively (Figure 17 C). GPT-4o performs qualitatively better, generating a more accurate expression at a significantly lower complexity than GPT-4. The MAE from both models is at least two to three times worse than the top-performing model identified by the BMS. BMS uses MCMC-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392. However, unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression. Our model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions. The optimal expression chosen is the best from 5 runs (1/153*5 or 1/765). The incremental suggestions from LLMs for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS. However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost. For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for five runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
@@
BMS samples expressions from a probability distribution. Therefore, running for a long time is expected to be characterized by equilibration, in which the expressions converge to a region in the stationary distribution of expressions, after which no significant improvement would be observed apart from continuing to explore the region of most likely expressions. Genetic algorithms like Eureqa [86] do not converge to a stationary distribution, nor does our approach. On the contrary, running a chatbot for a very long time leads to degradation of its answers as the context gets longer [63, 88], so, in principle, our method may exhibit similar degradation with very long runs. However, we expect this effect to be mitigated because we discard the majority of the context after each iteration, only passing the top and recent expressions.
Figure 18: Nikuradse models across the state-of-the-art symbolic regression programs. Here we compare our best suggested model from GPT-4o with other candidate models suggested from BMS, Eureqa, and EFS programs as reported in [40]. All three methods’ featured models are at different points on the Pareto Front; GPT-4o generated few equations in the regime where Eureqa / BMS found optimal equations.